Dynasty Corp. will pay a $3 dividend in one year. If investors expect that dividend to remain constant forever, and they require a 10% return on Dynasty stock, what is the stock worth? What is the stock worth if investors expect Dynasty’s dividends to grow at 3% per year?
Answer #1

a.Stock value=Annual dividend/Required return

=(3/0.1)=$30

b.Current price=D1/(Required return-Growth rate)

=3/(0.1-0.03)

=$42.86(Approx).